OpenStudy (anonymous):

The sides of a triangle are in the ratio of 2:3:4. If the measure of the shortest side of the triangle is 12 centimeters, what is the length of the longest side? 16 cm 18 cm 24 cm 36 cm

OpenStudy (lgbasallote):

12 corresponds to 2 yes?

OpenStudy (lgbasallote):

since 12 is the shortest side and in your ratio 2:3:4...2 is the shortest...so 12 corresponds to 2 right?

OpenStudy (anonymous):

24

OpenStudy (lgbasallote):

lol

OpenStudy (mertsj):

\[\frac{2}{12}=\frac{3}{x}=\frac{4}{y}\]

OpenStudy (lgbasallote):

or... \[\frac{12}{2} = \frac{x}{3} = \frac{y}{4}\] same thing really